OpenStudy (anonymous):

evaluate this integral ∫x-1/x dx

OpenStudy (lgbasallote):

\[\int \frac{x-1}{x}dx\] is that it?

OpenStudy (anonymous):

yup :)

OpenStudy (lgbasallote):

\[\int \frac{x-1}{x}dx \implies \int \frac{x}{x}dx - \frac{1}{x}dx \implies \int dx - \int \frac 1x dx\] does that help?

OpenStudy (lgbasallote):

maybe i should write it more beautifully \[\int dx - \int \frac{dx}{x}\] heh

OpenStudy (anonymous):

the answer will be -In|x| ?

OpenStudy (lgbasallote):

that's just \[-\int \frac{dx}{x}\]

OpenStudy (lgbasallote):

what about \[\int dx\]

OpenStudy (anonymous):

doesnt x/x cancel out ?

OpenStudy (lgbasallote):

it becomes 1

OpenStudy (lgbasallote):

do \[\int \frac xxdx \implies \int dx\]

OpenStudy (lgbasallote):

so*

OpenStudy (anonymous):

oh okay got it! thank you :)

OpenStudy (lgbasallote):

also dont forget +C

OpenStudy (lgbasallote):

since it's indefinite integral

OpenStudy (anonymous):

oh yeah